OpenStudy (jacobbirddog):

Plz Help!!! Fan And Medal! Solve the given number problem. The sum of three consecutive even integers is –72. What is the least of these numbers? A. –26 B. –22 C. –30 D. 24 (took a long time to type this -_-)

hartnn (hartnn):

you know what 'consecutive' mean ? (and what 'even' mean?)

OpenStudy (jacobbirddog):

consecutive = one after another even = 2,4,6,8,etc

OpenStudy (anonymous):

x +( x+2) + (x + 4) = 72 3x + 6 = 72 3x = 72 - 6 3x = 66 x = 22 Numbers are 22, 24 and 26

hartnn (hartnn):

good! so that will help you to make your equation if the least number is x then other 2 numbers are x+2 and x+4 got this ?

hartnn (hartnn):

hence your equation becomes x+ x+2 +x+4 = -72

OpenStudy (anonymous):

22,24,26 ....

hartnn (hartnn):

guys, its -72 there.

OpenStudy (jacobbirddog):

hmm. well im barely starting agebra so if i ask nooby questions sorry :P

hartnn (hartnn):

no problem! ask away :)

OpenStudy (anonymous):

Well then just change those numbers to negative and there's your answer

hartnn (hartnn):

not actually

OpenStudy (jacobbirddog):

so it would be -22 -24 -26 if i was to use the consecutive "rule" and i would use evens so that's why im saying -22 -24 -26?

OpenStudy (jacobbirddog):

ugh im a slow typer when it comes to these things... im a java-scripter not a expert on mathematics. lol.

hartnn (hartnn):

did u get how to get till here, x+ x+2 +x+4 = -72 ??

OpenStudy (jacobbirddog):

pardon?

OpenStudy (jacobbirddog):

did i get how to get to here... im not sure like i said im literally a noob it's my 2nd algebra lesson of the year and i can't finish the hw

hartnn (hartnn):

the 3 numbers are correct but we generally write in increasing order -26, -24, -22 now you can tell which one is the least :)

OpenStudy (jacobbirddog):

-26 would obviously be the least

hartnn (hartnn):

and that would be correct :)

OpenStudy (jacobbirddog):

so my answer would be: -26?

hartnn (hartnn):

YES

OpenStudy (jacobbirddog):

this is cool... i like algrbra :)

OpenStudy (jacobbirddog):

*algebra

hartnn (hartnn):

i'm glad you like it, you'll have fun then :)

OpenStudy (jacobbirddog):

yus thanks for all the help @hartnn

hartnn (hartnn):

most welcome ^_^